Lucy paid a total of $550 for a brooch, a necklace and a watch. She spent 3/5 money on the brooch. The necklace cost 3 times as much as the watch. How much did she pay for the necklace?
Found 2 solutions by ikleyn, math_tutor2020:
Answer by ikleyn(52790)

.
Lucy paid a total of $550 for a brooch, a necklace and a watch.
She spent 3/5 money on the brooch. The necklace cost 3 times as much as the watch.
How much did she pay for the necklace?
~~~~~~~~~~~~~

To simplify the problem, subtract the cost of the broch from $550.


The rest is 2/5 of $550, which is $220.


So, the necklace and the watch cost together $220, and the necklace costs 3 times as much as the watch.


From it, using your mental skills, you conclude that the necklace costs 3/4 of $220, or 3*55 = 165 dollars.    ANSWER




Alternatively, if you rely more to equations than on your mental skills,

you can write x + 3x = 220 and solve it for x, which is the price for the watch.


Doing this way, you get that the watch costs x= 55 dollars and the necklace costs 3x = 3*55 = 165 dollars, 
giving the same answer.

Solved.



Answer by math_tutor2020(3817)


w = cost of the watch in dollars
3w = cost of the necklace (3 times as much as the watch)

w+3w = 4w = combined cost of the watch and necklace

Lucy spent $550 for the brooch, necklace, and watch.
She spent 3/5 of that total on the brooch alone.
(3/5)*550 = 330 dollars was spent on the brooch alone

The amount left over is 550-330 = 220
Set this equal to the 4w mentioned and solve for w.

4w = 220
w = 220/4
w = 55
The watch is $55

Necklace = 3w = 3*55 = 165

Answer: The necklace is $165